OpenStudy (anonymous):

What is 4.6 × 10^8 cm/min in meters per hour? A. 2.76 × 10^8 m/h B. 2.76 × 10^10 m/h C. 7.67 × 10^6 m/h D. 7.67 × 10^8 m/h

OpenStudy (anonymous):

1h=? min

OpenStudy (anonymous):

?

OpenStudy (anonymous):

60.

OpenStudy (anonymous):

ok, 1m=?cm?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

\[4.6 × 10^8 \frac{ cm }{ \min }\times \frac{ 60\min }{ 1h }\times \frac{ 1m }{ 100cm }=\]

OpenStudy (anonymous):

can you do it ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

we have 1m=100cm and 1h=60min

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we need this quantity in m/h, so we cancel first min and replace it with h

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[4.6 × 10^8 \frac{ cm }{ \min }\times \frac{ 60\min }{ 1h }\]